OpenStudy (anonymous):

A 50 kg block on the floor. The coefficients of static and kinetic friction are 0.7 and 0.5 respectively. (a) What is the minimum force needed to start the block moving? (b) If the same force continues to push the block after it starts moving, what is the acceleration of the block?

OpenStudy (anonymous):

for minimum force ...use static friction

OpenStudy (anonymous):

and after it starts moving ...use kinetic friction for friction force.

OpenStudy (anonymous):

ATTEMPT TO THE SOLUTION: a) a=0 let f= friction \[\Sigma F _{x} = ma\] F-f = 0 \[\Sigma F _{y} = 0\] N=W \[F= muW\] \[F= (0.7)(50)(9.8)\] F= 343 N b.) a=? \[SigmaF _{x}= ma\] F-f=ma \[SigmaF _{y} = 0\] N=W \[a= \frac{ F-f }{ m }\] \[a= \frac{ (343)- (0.5) }{ 50 }\] a= 6.85 m/s^2 is this correct?

OpenStudy (anonymous):

a= (343-0.5)/50 ?

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

how did u get this ?

OpenStudy (anonymous):

isn't (343- 0.5*m*g)/50 =a

OpenStudy (anonymous):